An aluminium conductor, 1 km long has a diameter of 0.5 cm. Determine the resistance at 35 °C.



1
Expert's answer
2021-03-30T19:54:30-0400
"R=\rho\frac{l}{A}\\R=(2.7\cdot10^{-8})\frac{1000}{0.25\pi(0.005)^2}=1.375\Omega"